OpenStudy (anonymous):

how do you write 4/5 x 10^11 in standard form?

sam (.sam.):

8x10^10

OpenStudy (anonymous):

\[\frac{4}{5}\cdot 10^{11}=0.8\cdot 10\cdot 10^{10}=8\cdot 10^{10}\]

OpenStudy (radar):

80,000,000,000

OpenStudy (anonymous):

thanks radar that's what i thought!!